# Theory Uncertainties in the Extraction of $\alpha_{s}$ from Drell-Yan at Small Transverse Momentum

### v1

@@ 1 Introduction
The strong coupling constant $\alpha_{s}$, which parameterizes the strength of the strong interactions, is the least well known among the fundamental forces of nature. Its precise determination has thus been of great interest with many complementary determinations from different sources being pursued over the years [1]. The current world average of $\alpha_{s}(m_{Z})=0.1180(9)$ [2] has a relative uncertainty of $0.8$%, with the most precise result stemming from the average of lattice QCD determinations $\alpha_{s}(m_{Z})=0.1183(7)$ ($0.6\%$[3].
One of the currently most precise determinations from experimental measurements at colliders comes from fits to hadronic event shapes in $e^{+}e^{-}$ collisions [4], most notably the thrust distribution [5, 6] yielding $\alpha_{s}(m_{Z})=0.1136(12)$ [7] with a precision of about 1%, which however has shown a persistent tension with the world average, despite much theoretical scrutiny [8, 9, 10, 11, 12, 13, 14, 7, 15]. An independent determination of similarly high precision at or below the one-percent level from collider measurements is thus of great importance. A promising candidate for such a high-precision determination is the transverse-momentum ($q_{T}$) spectrum in Drell-Yan production, which has been measured to subpercent precision at the LHC [16, 17, 18, 19, 20, 21, 22, 23, 24] and for which theoretical predictions have been reaching the same N3LL and (approximate) N4LL resummation as for $e^{+}e^{-}$ event shapes resulting in a perturbative precision at the few-percent level [25, 26, 27, 28, 29, 30, 31, 32].
@@
Another important source of uncertainty is due to the limited knowledge of parton distribution functions (PDFs). Whilst the propagation of their uncertainties is in principle a more straightforward exercise, in practice it faces some nontrivial caveats. We therefore do not consider them here but will address them in a dedicated paper [35].
@@ 2 Theory Uncertainties and Correlations for Differential Spectra
We will generically use $f_{n}(x)$ to denote the expansion coefficients and specifically $\hat{f}_{n}(x)$ to denote their true values in nature. Calculating the first few true coefficients, we obtain a theory prediction for $f(x)$ at leading order (LO), next-to-leading order (NLO), next-to-next-to-leading order (NNLO), and so on,
@@ 2.2 Limitations of scale variations
NLO: $\tilde{f}(x,\tilde{\alpha})$ $=\hat{\tilde{f}}_{0}(x)+\hat{\tilde{f}}_{1}(x)\,\tilde{\alpha}$
@@
NNLO: $\Delta f(x,\alpha)$ =$=\bigl{\{}2b_{0}[\hat{f}_{2}(x)-b_{0}\hat{f}_{1}(x)]+b_{1}\hat{f}_{1}(x)\bigr{\}}\alpha^{3}+\mathcal{O}(\alpha^{4})\,.$ (2.7)
@@
$a$ 00 $-1$ 00
$b$ $+1$ $+1$ $+1$
$c$ $+1$ $-1$ $-1$
Figure -959: Interpreting a differential spectrum by scanning scale variations. Left: The theory prediction (black line) with a theory uncertainty band (blue) from the envelope of scale variations compared to measurements (red) at three points $x_{1,2,3}$. Middle: The normalized uncertainty band (blue) with three possible scale variations (orange, green, violet) filling out the band. Right: The assumed correlations $\rho_{ij}$ between the uncertainties at the points $x_{i}$ and $x_{j}$ resulting from a given scale variation. Note that for $\rho_{ij}=0$ the uncertainty at either $x_{i}$ or $x_{j}$ must vanish.
@@
Instead, this scanning over enveloped variations amounts to trying out various correlation models for the same total uncertainty band, as illustrated in figure -959. We might then wonder what we can learn from this. First, it is important to realize that none of the trial variations actually provides a realistic correlation model, because as seen in figure -959, they are by construction always fully correlated, only allowing correlations $\rho_{ij}=\pm 1$ (or $\rho_{ij}=0$ in case $\Delta f(x_{i,j})=0$). This is clearly problematic when one is sensitive to shape effects, where as discussed in section 2.1, the precise correlation near $\rho=1$ completely determines the resulting shape uncertainty and thus the impact on the parameter of interest. Therefore any given variation with its implied correlation model can strongly underestimate (or overestimate) the resulting theory uncertainty on the parameter of interest. One might be “lucky” such that one trial variation sufficiently closely mimics the shape effect of the parameter of interest so as to yield a “conservative” uncertainty and avoid underestimation. However, this might just as well be very “unlucky” and yield a substantial overestimation.
@@ 2.3 Theory nuisance parameters
Here, the (considered to be) unknown series coefficients $f_{n}(x,\theta_{n})$ are parameterized in terms of theory nuisance parameters $\theta_{n}$, where for simplicity we let $\theta_{n}\equiv\{\theta_{n,i}\}$ denote the set of all scalar parameters $\theta_{n,i}$.
@@
[…]
@@ 3.1 Theory requirements
Beyond these leading contributions, achieving a (sub)percent-level theoretical description requires incorporating various subleading effects. One such effect is the nonsingular contribution in Eq. (3.1). Similarly, quark mass corrections scaling as $m_{q}^{2}/q_{T}^{2}$, and QED and electroweak effects become relevant at this level of precision. Neglecting these subleading effects can cause a substantial bias in the extracted value of $\alpha_{s}(m_{Z})$, so it is important to include them in a fit to real data. However, once they are included, we do not expect them to have a major influence on the uncertainty in $\alpha_{s}(m_{Z})$. We can therefore consistently neglect them in our study as further discussed in section 3.4.
@@ 3.2 Resummation of perturbative contributions at small $q_{T}$
×$\quad\times\int\!\!\frac{\mathrm{d}^{2}{\vec{b}}_{T}}{(2\pi)^{2}}\,e^{\mathrm{i}{\vec{b}}_{T}\cdot{\vec{q}}_{T}}\,\tilde{B}_{a}(x_{a},b_{T},\mu,\nu/Q)\,\tilde{B}_{b}(x_{b},b_{T},\mu,\nu/Q)\tilde{S}(b_{T},\mu,\nu)\,.$
@@ 3.2.1 Resummation
In the perturbative region $q_{T}\sim 1/b_{T}\gg\Lambda_{\mathrm{QCD}}$, this boundary condition can be computed in terms of standard PDFs at some scale $\mu_{f}$,
$\tilde{b}_{i,n}(x)$ $=\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\tilde{I}_{ij,n}(z,\mu_{f})\,f_{j}\Bigl{(}\frac{x}{z},\mu_{f}\Bigr{)}\,,$ (3.7)
@@
$\mu_{H}$ $=Q\,,\qquad$
$\mu_{B}$ $=Q\,f_{\rm run}\biggl{(}\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\mu_{B}^{\rm min}\Bigr{)}\biggr{)}\,,\qquad$ $\nu_{B}$ $=Q\,,$
$\mu_{S}$ $=Q\,f_{\rm run}\biggl{(}\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\mu_{S}^{\rm min}\Bigr{)}\biggr{)}\,,\qquad$ $\nu_{S}$ $=Q\,f_{\rm run}\biggl{(}\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\nu_{S}^{\rm min}\Bigr{)}\biggr{)}\,,$
$\mu_{0}$ $=\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\mu_{0}^{\rm min}\Bigr{)}\,,\qquad$ $\mu_{f}$ $=Q\,f_{\rm run}\biggl{(}\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\mu_{f}^{\rm min}\Bigr{)}\biggr{)}\,,$ (3.9)
@@ 3.2.2 Scale variations
$\mu_{B}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\mu_{B}}}\,\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\frac{\mu_{B}^{\rm min}}{2^{w_{\rm FO}}f_{\rm vary}^{v_{\mu_{B}}}}\Bigr{)}\,,\qquad$ $\nu_{B}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\nu_{B}}}\,Q\,,$
$\mu_{S}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\mu_{S}}}\,\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\frac{\mu_{S}^{\rm min}}{2^{w_{\rm FO}}f_{\rm vary}^{v_{\mu_{S}}}}\Bigr{)}\,,\qquad$ $\nu_{S}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\nu_{S}}}\,\mu_{*}\Bigl{(}\frac{b_{0}}{b_{T}},\frac{\nu_{S}^{\rm min}}{2^{w_{\rm FO}}f_{\rm vary}^{v_{\nu_{S}}}}\Bigr{)}\,,$
@@ 3.2.3 TNPs for $q_{T}$ resummation
The generalized order counting for resummation including TNPs is denoted as Nm+kLL and defined in analogy to Nm+kLO in section 2.3 as follows: A given perturbative resummation order, NnLL, is uniquely and well defined by including all underlying perturbative series in the RGE to a certain order in $\alpha_{s}$, namely the boundary conditions to $(n-1)$ loops, the noncusp anomalous dimensions to $n$ loops, and the cusp anomalous dimensions and QCD beta functions to $n+1$ loops. At Nm+kLL we use the complete resummation structure at NnLL with $n=m+k$ with the true values for all coefficients relevant up to NmLL and the TNP parameterization for the last $k$ coefficients of each series. For example, Eq. (3.15) at N3+1LL takes the form
@@
$\tilde{I}_{ij}(z,\theta_{n}^{B_{ij}})=\frac{3}{2}\theta_{n}^{B_{ij}}\hat{\tilde{I}}_{ij,n}(z)\,,$ (3.21)
@@ 3.3.1 General discussion
At large $b_{T}\sim 1/\Lambda_{\mathrm{QCD}}$, $\tilde{f}^{\mathrm{np}}_{i}(x,b_{T})$ and $\tilde{\gamma}_{\zeta}^{\mathrm{np}}(b_{T})$ are genuine nonperturbative functions of $b_{T}$. However, the relevant region of interest for our purposes is the spectrum for $q_{T}\gtrsim 1\,\mathrm{GeV}$, which is determined by much smaller $b_{T}$, namely where $1/b_{T}\sim q_{T}\gg\Lambda_{\mathrm{QCD}}$ is still in the perturbative domain. In this region, nonperturbative effects are not negligible, but we can systematically expand them in an operator product expansion (OPE) in powers of $\Lambda_{\mathrm{QCD}}b_{T}\ll 1$ [88, 69, 70, 89, 90, 91, 92]. The OPE of the final TMD PDF entering the factorized cross section takes the form [92]
@@ 3.3.2 Nonperturbative parameterization
=$=\begin{cases}-\lambda_{2}\,b_{T}^{2}-\lambda_{4}\,b_{T}^{4}+\mathcal{O}(b_{T}^{6})\qquad&1/b_{T}\gg\Lambda_{\mathrm{QCD}}\\ -\lambda_{\infty}&1/b_{T}\ll\Lambda_{\mathrm{QCD}}\,,\end{cases}$
@@
=$=\begin{cases}-\Lambda_{2}\,b_{T}^{2}-\Lambda_{4}\,b_{T}^{4}+\mathcal{O}(b_{T}^{6})\qquad&1/b_{T}\gg\Lambda_{\mathrm{QCD}}\\ -\Lambda_{\infty}\,b_{T}&1/b_{T}\ll\Lambda_{\mathrm{QCD}}\,,\end{cases}$
@@
We note that the above model for the TMD PDF does not yet include flavor or $x$ dependence, which would be required for a more complete treatment. As we explain in section 3.4, this simplified setup is sufficient for Asimov fits since we can use the same model in both pseudodata and fitted theory model. However, it is likely not sufficient for a fit to real data. As shown in Ref. [32], the full flavor and $x$ dependence of the nonperturbative TMD model can be approximated by an effective flavor-averaged model that only depends on the rapidity $Y$ and a given resonant boson type and center-of-mass energy. However, this also implies that once several different rapidity bins are considered, the effective TMD parameters must at least depend on rapidity. The fits to the $q_{T}$ spectrum performed in Ref. [57] found indeed evidence for the presence of a nontrivial effective rapidity dependence of $\Lambda_{2}$.
For the purpose of our Asimov fits, we also need some representative values for the nonperturbative parameters to be used as true values in our Asimov pseudodata. As already mentioned above, the nonperturbative CS kernel can be constrained by lattice QCD data [84, 85, 83]. This requires properly accounting for quark flavor thresholds and quark mass effects as will be discussed in Ref. [87]. From there, we obtain some representative values for the CS kernel parameters of our model,
@@
with the corresponding correlation matrix
@@
For the TMD parameters, due to a lack of robust constraints, we choose a generic size of $(0.5\,\mathrm{GeV})^{n}$ as representative central values. Specifically, we take

### v2

@@ 1 Introduction
The strong coupling constant $\alpha_{s}$, which parameterizes the strength of the strong interactions, is the least well known among the fundamental forces of nature. Its precise determination has thus been of great interest with many complementary determinations from different sources being pursued over the years [1]. The current world average of $\alpha_{s}(m_{Z})=0.1180(9)$ [2] has a relative uncertainty of 0.80.8%, with the most precise result stemming from the average of lattice QCD determinations $\alpha_{s}(m_{Z})=0.1183(7)$ ($0.6\%$[3].
One of the currently most precise determinations from experimental measurements at colliders comes from fits to hadronic event shapes in $e^{+}e^{-}$ collisions [4], most notably a determination from the thrust distribution [5, 6, 7] yielding $\alpha_{s}(m_{Z})=0.1136(12)$ with a precision of about 1%, which however has shown a persistent tension with the world average, despite much theoretical scrutiny [8, 9, 10, 11, 12, 13, 14, 7, 15]. An independent determination of similarly high precision at or below the one-percent level from collider measurements is thus of great importance. A promising candidate for such a high-precision determination is the transverse-momentum ($q_{T}$) spectrum in Drell-Yan production, which has been measured to subpercent precision at the LHC [16, 17, 18, 19, 20, 21, 22, 23, 24] and for which theoretical predictions have been reaching the same N3LL and (approximate) N4LL resummation as for $e^{+}e^{-}$ event shapes resulting in a perturbative precision at the few-percent level [25, 26, 27, 28, 29, 30, 31, 32].
@@
Another important source of uncertainty is due to the limited knowledge of parton distribution functions (PDFs), which are required inputs for such hadron collider determinations of $\alpha_{s}(m_{Z})$. The PDFs themselves have a dependence on $\alpha_{s}$ [35, 36, 37, 38, 39], and this non-trivial PDF-$\alpha_{s}$ correlation must be propagated faithfully to the extracted output $\alpha_{s}(m_{Z})$. Whilst the propagation of these uncertainties is in principle a more straightforward exercise, in practice it faces some nontrivial caveats. We therefore do not consider them here but will address them in a dedicated paper [40].
@@ 2 Theory Uncertainties and Correlations for Differential Spectra
We will generically use $f_{n}(x)$ to denote the expansion coefficients and specifically $\hat{f}_{n}(x)$ to denote their true values. Calculating the first few true coefficients, we obtain a theory prediction for $f(x)$ at leading order (LO), next-to-leading order (NLO), next-to-next-to-leading order (NNLO), and so on,
@@ 2.2 Limitations of scale variations
[…]
@@
NNLO: $\Delta f(x,\alpha)$ $=\bigl\{2b_{0}[\hat{f}_{2}(x)-b_{0}\hat{f}_{1}(x)]+b_{1}\hat{f}_{1}(x)\bigr\}\alpha^{3}+\mathcal{O}(\alpha^{4})\,.$ (2.7)
@@
$a$ 0 1-1 0
$b$ +1+1 +1+1 +1+1
$c$ +1+1 1-1 1-1
Figure 1: Interpreting a differential spectrum by scanning scale variations. Left: The theory prediction (black line) with a theory uncertainty band (blue) from the envelope of scale variations compared to measurements (red) at three points $x_{1,2,3}$. Middle: The normalized uncertainty band (blue) with three possible scale variations (orange, green, violet) filling out the band. Right: The assumed correlations $\rho_{ij}$ between the uncertainties at the points $x_{i}$ and $x_{j}$ resulting from a given scale variation. Note that for $\rho_{ij}=0$ the uncertainty at either $x_{i}$ or $x_{j}$ must vanish.
@@
Instead, this scanning over enveloped variations amounts to trying out various correlation models for the same total uncertainty band, as illustrated in figure 1. We might then wonder what we can learn from this. First, it is important to realize that none of the trial variations actually provides a realistic correlation model, because as seen in figure 1, they are by construction always fully correlated, only allowing correlations $\rho_{ij}=\pm 1$ (or $\rho_{ij}=0$ in case $\Delta f(x_{i,j})=0$). This is clearly problematic when one is sensitive to shape effects, where as discussed in section 2.1, the precise correlation near $\rho=1$ completely determines the resulting shape uncertainty and thus the impact on the parameter of interest. Therefore any given variation with its implied correlation model can strongly underestimate (or overestimate) the resulting theory uncertainty on the parameter of interest. One might be “lucky” such that one trial variation sufficiently closely mimics the shape effect of the parameter of interest so as to yield a “conservative” uncertainty and avoid underestimation. However, this might just as well be very “unlucky” and yield a substantial overestimation.
@@ 2.3 Theory nuisance parameters
Here, the (considered to be) unknown series coefficients $f_{n}(x,\theta_{n})$ are parameterized in terms of theory nuisance parameters $\theta_{n}$, where for simplicity we let $\theta_{n}\equiv\{\theta_{n,i}\}$ denote an appropriate set of parameters $\theta_{n,i}$ that are by construction scalars [34].
@@
which can come from theory considerations or experimental constraints or both. Without loss of generality we will assume that, prior to applying any experimental constraints, the $\theta_{n}$ are chosen such that based on theory information only we have $u_{n}=0$ and $\Delta u_{n}\equiv\Delta\theta_{n}=1$. [^2] Starting from a theory constraint $u_{n}\pm\Delta u_{n}$ for $\theta_{n}$, we can always reparameterize it as $\theta_{n}^{\prime}=(\theta_{n}-u_{n})/\Delta u_{n}$, such that the corresponding constraint on $\theta_{n}^{\prime}$ is $0\pm 1$.
@@ 3.1 Theory requirements
Beyond these leading contributions, achieving a (sub)percent-level theoretical description requires incorporating various subleading effects. One such effect is the nonsingular contribution in Eq. (3.1). Similarly, quark mass corrections scaling as $m_{q}^{2}/q_{T}^{2}$, and QED and electroweak effects become relevant at this level of precision (see e.g. Refs. [65, 66, 67, 68, 69, 70]). Neglecting these subleading effects can cause a substantial bias in the extracted value of $\alpha_{s}(m_{Z})$, so it is important to include them in a fit to real data. However, once they are included, we do not expect them to have a major influence on the uncertainty in $\alpha_{s}(m_{Z})$. We can therefore consistently neglect them in our study as further discussed in section 3.4.
@@ 3.2 Resummation of perturbative contributions at small $q_{T}$
$\quad\times\int\!\!\frac{\mathrm{d}^{2}{\vec{b}}_{T}}{(2\pi)^{2}}\,e^{\mathrm{i}{\vec{b}}_{T}\cdot{\vec{q}}_{T}}\,\tilde{B}_{a}(x_{a},b_{T},\mu,\nu/Q)\,\tilde{B}_{b}(x_{b},b_{T},\mu,\nu/Q)\tilde{S}(b_{T},\mu,\nu)\,.$
@@ 3.2.1 Resummation
In the perturbative region $q_{T}\sim 1/b_{T}\gg\Lambda_{\mathrm{QCD}}$, this boundary condition is computed in terms of standard PDFs at the canonical scale $\mu_{f}^{\rm can}=b_{0}/b_{T}$,
$\tilde{b}_{i,n}(x)$ $=\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\tilde{I}_{ij,n}(z)\,f_{j}\Bigl(\frac{x}{z},\mu_{f}^{\rm can}\Bigr)\,$, (3.7)
@@
$\mu_{H}$ $=Q\,$$\qquad$
$\mu_{B}$ $=Q\,f_{\rm run}\biggl(\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\mu_{B}^{\rm min}\Bigr)\biggr)\,$$\qquad$ $\nu_{B}$ $=Q\,$,
$\mu_{S}$ $=Q\,f_{\rm run}\biggl(\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\mu_{S}^{\rm min}\Bigr)\biggr)\,$$\qquad$ $\nu_{S}$ $=Q\,f_{\rm run}\biggl(\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\nu_{S}^{\rm min}\Bigr)\biggr)\,$,
$\mu_{0}$ $=\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\mu_{0}^{\rm min}\Bigr)\,$$\qquad$ $\mu_{f}$ $=Q\,f_{\rm run}\biggl(\frac{q_{T}}{Q},\frac{1}{Q}\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\mu_{f}^{\rm min}\Bigr)\biggr)\,$, (3.9)
@@ 3.2.2 Scale variations
$\mu_{B}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\mu_{B}}}\,\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\frac{\mu_{B}^{\rm min}}{2^{w_{\rm FO}}f_{\rm vary}^{v_{\mu_{B}}}}\Bigr)\,$$\qquad$ $\nu_{B}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\nu_{B}}}\,Q\,$,
$\mu_{S}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\mu_{S}}}\,\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\frac{\mu_{S}^{\rm min}}{2^{w_{\rm FO}}f_{\rm vary}^{v_{\mu_{S}}}}\Bigr)\,$$\qquad$ $\nu_{S}$ $=2^{w_{\rm FO}}\,f_{\rm vary}^{v_{\nu_{S}}}\,\mu_{*}\Bigl(\frac{b_{0}}{b_{T}},\frac{\nu_{S}^{\rm min}}{2^{w_{\rm FO}}f_{\rm vary}^{v_{\nu_{S}}}}\Bigr)\,$,
@@ 3.2.3 TNPs for $q_{T}$ resummation
The generalized order counting for resummation including TNPs is denoted as Nm+kLL and defined in analogy to Nm+kLO in section 2.3 as follows: A given perturbative resummation order, NnLL, is uniquely and well defined by including all underlying perturbative series in the RGE to a certain order in $\alpha_{s}$, namely the boundary conditions to (n1)(n-1) loops, the noncusp anomalous dimensions to $n$ loops, and the cusp anomalous dimensions and QCD beta functions to $n+1$ loops. At Nm+kLL we use the complete resummation structure at NnLL with $n=m+k$ with the true values for all coefficients relevant up to NmLL and the TNP parameterization for the last $k$ coefficients of each series. For example, Eq. (3.15) at N3+1LL takes the form
@@
$\tilde{I}_{ij,n}(z,\theta_{n}^{B_{ij}})=\frac{3}{2}\,\theta_{n}^{B_{ij}}\,\hat{\tilde{I}}_{ij,n}(z)\,$, (3.21)
@@ 3.3.1 General discussion
At large $b_{T}\sim 1/\Lambda_{\mathrm{QCD}}$, $\tilde{f}^{\mathrm{np}}_{i}(x,b_{T})$ and $\tilde{\gamma}_{\zeta}^{\mathrm{np}}(b_{T})$ are genuine nonperturbative functions of $b_{T}$. However, the relevant region of interest for our purposes is the spectrum for $q_{T}\gtrsim 1\,\mathrm{GeV}$, which is determined by much smaller $b_{T}$, namely where $1/b_{T}\sim q_{T}\gg\Lambda_{\mathrm{QCD}}$ is still in the perturbative domain. In this region, nonperturbative effects are not negligible, but we can systematically expand them in an operator product expansion (OPE) in powers of $\Lambda_{\mathrm{QCD}}b_{T}\ll 1$ [93, 75, 76, 94, 95, 96, 97] (see also Ref. [98]). The OPE of the final TMD PDF entering the factorized cross section takes the form [97]
@@ 3.3.2 Nonperturbative parameterization
[…]
@@
[…]
@@
We note that the above model for the TMD PDF does not yet include flavor or $x$ dependence, which would be required for a more complete treatment. As we explain in section 3.4, this simplified setup is sufficient for Asimov fits since we can use the same model in both pseudodata and fitted theory model. However, it is likely not sufficient for a fit to real data, which in general requires flavor and $x$ dependence [99, 31, 100]. As shown in Ref. [32], the full flavor and $x$ dependence of the nonperturbative TMD model can be approximated by an effective flavor-averaged model that only depends on the rapidity $Y$ and a given resonant boson type and center-of-mass energy. However, this also implies that once several different rapidity bins are considered, the effective TMD parameters must at least depend on rapidity. The fits to the $q_{T}$ spectrum performed in Ref. [62] found indeed evidence for the presence of a nontrivial effective rapidity dependence of $\Lambda_{2}$.
For the purpose of our Asimov fits, we also need some representative values for the nonperturbative parameters to be used as true values in our Asimov pseudodata. As already mentioned above, the nonperturbative CS kernel can be constrained by lattice QCD data [90, 91, 89]. This requires properly accounting for quark flavor thresholds and quark mass effects as will be discussed in Ref. [66]. Using Ref. [66] to fit our nonperturbative parameterization of the CS kernel to the lattice QCD data, we obtain the following results,
@@
with the resulting correlation matrix
@@
We will use these as representative values for the CS kernel parameters of our model. For the TMD parameters, due to a lack of robust constraints, we choose a generic size of $(0.5\,\mathrm{GeV})^{n}$ as representative central values. Specifically, we take